A fairground ride spins its occupants inside a flying saucer-shaped container. If the horizontal circular path the riders follow has a 6.3 m radius, if the riders is subjected to a centripetal acceleration 1.3 times that due to gravity, what is the spinning speed in unit of revolutions per minute? Use g = 9.8 m/s2.
Answer #1

Solution)

Given,

Radius, r=6.3 m

Centripetal acceleration, a=1.3*g

We know, centripetal acceleration, a=rw^2

So, 6.3*w^2=1.3*9.8

Angular velocity, w= 1.422 rad/s

We also know, Time period, T for one revolution=2pi/w=2*pi/1.422= 4.41 sec

Now, spinning speed in rev/min=60/4.41= 13.58 rev/min (Ans)

=========